Which of the following numbers is equal to one-fourth of its own reciprocal?
a) 1/8 b) 1/4 c) 1/2 d) 2 e) 4

Answer by Theo(13342)

looks like number C.

1/2 is equal to 1/4 of 2.

if you divide 2 by 1/2 you will get 4.

1/2 / 2 = 1/2 * 1/2 = 1/4